OpenStudy (anonymous):

Related Rates: I know I have to do derivatives but I'm stuck on how to get from being given dv/dt to finding area. Also confused about finding radius when not given dr/dt only the height A chemical tank is leaking a toxic substance at a rate of 100cm^3/minute. Slick is semi-circ. in shape and has a thicknesss of .5cm a) After 1 hour how big an area does the slick cover? b) At 1 hour, at what rate is the radius of the slick changing?

OpenStudy (amistre64):

i think if we just half the height, we can create a full circle model

OpenStudy (amistre64):

V = pir^2 h/2 V' = 2pi r r' h/2 + pi r h'/2 ; but h' = 0 since the height is not changing V' = pi r r' h 100 = .25 pi r r' 400/pi = r r' ------------------------------------ area covered is just the base of the circle regardless of height A = pi r^2 A' = 2pi r r' ... per minute, we already have determined the r r' = 400/pi A' = 2pi(400/pi) = 800 cm^2 per minute

OpenStudy (amistre64):

use 60 minutes to determine the area after an hour use the Area at an hour to determine the radius, and solve for r' in the V' setup

OpenStudy (amistre64):

might have some errors in calculations ... but the concept should be sound

OpenStudy (amistre64):

im thinking i mentally halved the height to begin with and it should be: 100 = .5 pi r r' 200/pi = r r'